Sanjana loves to read books(B) and likes to watch movies (M) as well. She derives double the satisfaction from reading a book than watching a movie. Her utility function is as follows








U(B, M) = 2B + M








Where B and M are the number of books and movies respectively.








a. Explain the shape of the indifference curve using diagrams. [3 MARKS, Word limit: 200] b. The income available with her is INR60. And price of movie ticket is INR 10 and that of a book is INR 15. Will she spend her entire budget and why? What will be her consumption, explain why?

1
Expert's answer
2021-12-06T09:36:50-0500

"\left\{\begin{matrix}
 15B+10M\leqslant 60, \\ 
 B=2M; 
\end{matrix}\right."


"\left\{\begin{matrix}
 M\leqslant 1.5, \\ 
B=3; 
 
\end{matrix}\right."


"\left\{\begin{matrix}
 M= 1, \\ 
B=3; 
 
\end{matrix}\right."


she will not spend her entire budget,

her consumption will be

"60-1\cdot 10-3\cdot 15=5" INR.